1).Β Two strings support a load of weight 7.5N, One string has a tension 𝑇1 and is at an angle 50Β° to the horizontal. The other string has a tension 𝑇2 and is at angle 40Β° to the horizontal. The object is in equilibrium. Determine the values of 𝑇1 and 𝑇2 by using a vector triangle or by resolving forces. [4marks]


2). A cylindrical disc has diameter 28mm and thickness 12mm. the material of the disc has density 6.8 Γ— 103kgmβˆ’3 . Calculate, to two significant figures, the weight of the disc.


1
Expert's answer
2021-10-01T12:22:33-0400

1) Write the equilibrium of forces along y- and x-axes:


"y: T_1\sin50°+T_2\sin40°-W=0,\\
x: -T_1\cos50°+T_2\cos40°=0.\\\space\\
T_1=T_2\frac{\cos40°}{\cos50°}=1.19T_2.\\\space\\
(1.19T_2)\sin50°+T_2\sin40°=W,\\\space\\
T_2=\frac{W}{1.19\sin50°+\sin40°}=4.8\text{ N},\\\space\\
T_1=1.19T_2=5.7\text{ N}."

2) The weight is


"W=mg=\rho Vg=\rho \frac{\pi d^2}{4}tg=0.49\text{ N}."